OpenStudy (anonymous):

If -2 is a zero of the polynomial root 2 (x+p) and is also the zero of the polynomial px^2+kx+2 root 2then find the value of k

OpenStudy (loser66):

1) put -2 into \(\sqrt 2 (x+p) \) =0 to find p 2) replace p into px^2 +kx +2 3) solve p (-2)^2 +k(-2) +2 =0 for k